a) Suppose the demand for a product is given by p=d(q)=-3q²+6q+436  where  q is the quantity and p is the price in Ghana Cedis. Using calculus, estimate the effect of an increase in price from GH¢196.00 to GH¢250.00 on the welfare of the consumer.Using the concept of own elasticity of demand, decide whether the producer should accept this price increase in order to increase revenue.

1
Expert's answer
2021-11-07T19:47:02-0500
"-3q^2+6q+436=196"

"q=10"

"pq=1960"

"-3q^2+6q+436=250"

"q=9"

"pq=2250"

"\varDelta pq=1960-2250=-290"

Consumer welfare decreased by 290 Ghanaian cedis. The manufacturer accepts a price increase to increase revenue.